OpenStudy (anonymous):

Use the function: 7e^3^x = 312 The approximate solution would be: Select one: a. 0 b. 1.3 c. 1.99 d. 2.4

OpenStudy (anonymous):

@Michele_Laino i wanna say 0 because http://www.wolframalpha.com/input/?i=7e%5E3%5Ex+%3D+312

OpenStudy (anonymous):

or maybe C

OpenStudy (michele_laino):

I think that we can divide both sides by \(7\), so we can write: \[\Huge {e^{{3^x}}} \cong 45\]

OpenStudy (anonymous):

oh okay

OpenStudy (michele_laino):

subsequently, we can take the natural logaritm, like this: \[\Huge {3^x} \cong \ln 45\]

OpenStudy (michele_laino):

and finally, we can take the logarithm with base 3: \[\Huge x \cong {\log _3}\left( {\ln 46} \right)\] what do you think about procedure? @Jhannybean

OpenStudy (jhannybean):

Seems good to me.

OpenStudy (michele_laino):

thanks!! :) @Jhannybean

OpenStudy (anonymous):

Okay so do we solve that or do nother step?

OpenStudy (michele_laino):

it is simplewe have to evaluate this quantity: \[\Huge x \cong {\log _3}\left( {\ln 45} \right) = ...?\]

OpenStudy (anonymous):

i dont know how

OpenStudy (michele_laino):

first step: using a scientific calculator, we can write this: \[\Huge \ln 45 \cong 3.81\]

OpenStudy (michele_laino):

second step: In order to evaluate the \(\log_3\) we have to change the basis of the logaritm, namely from 3 to 10, using this identity \[\huge {\log _3}\left( {\ln 45} \right) = \frac{{{{\log }_{10}}\left( {\ln 45} \right)}}{{{{\log }_{10}}3}} = ...?\]

OpenStudy (anonymous):

2.69

OpenStudy (michele_laino):

more steps: \[\huge \begin{gathered} {\log _3}\left( {\ln 45} \right) = \frac{{{{\log }_{10}}\left( {\ln 45} \right)}}{{{{\log }_{10}}3}} = \hfill \\ \hfill \\ = \frac{{{{\log }_{10}}3.81}}{{{{\log }_{10}}3}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

I got 1.22

OpenStudy (anonymous):

thats not an answer choice though?

OpenStudy (michele_laino):

I think that we can pick option B

OpenStudy (anonymous):

okay great Use the following equation: log2x + log2(x – 7) = 3 The third step would be to: Select one: a. Graph the function. b. Take the log of each side. c. Square the expression. d. Factor the resulting quadratic equation.

OpenStudy (michele_laino):

here we have arrived to this result: \[\huge {\log _2}x + {\log _2}\left( {x - 7} \right) = {\log _2}8\] right?

OpenStudy (anonymous):

right

OpenStudy (michele_laino):

now I can apply this property of logarithms: \(\log m +\log n=\log (m \cdot n)\) so I can write this: \[\huge {\log _2}\left\{ {x \cdot \left( {x + 7} \right)} \right\} = {\log _2}8\]

OpenStudy (anonymous):

oh okay

OpenStudy (michele_laino):

therefore, I equate both numbers of such logarithms, so I get: \[\huge \begin{gathered} x \cdot \left( {x + 7} \right) = 8 \hfill \\ \hfill \\ x \cdot \left( {x + 7} \right) - 8 = 0 \hfill \\ \end{gathered} \] and the last expression is a quadratic equation. SO, what is the right option?

OpenStudy (michele_laino):

oops... So*

OpenStudy (anonymous):

B?

OpenStudy (michele_laino):

no, since we have not the logaritms now, please we have a quadratic equation which we have to solve

OpenStudy (anonymous):

D

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (michele_laino):

:)